Fill in the truth table for each of the following logic circuits ABCF A B CIF Á DDF
Answer #1

The three circuits with the state at their input and output are as follows
U1 AB NAND2 A.B U11 F=(A.B). (B+C) U2 AND2 B+C B+C NOR2 U6 A.B D -AB U10 U10 AND2 - F=A.BOC F=A.B® © Vē XOR2 Da XOR2U3 AOC AOC U8 XOR2 F= AOC + BOC U7 Do OR2 XNOR2 BOC Вос

please consider the XNOR sign. I tried my best to make it.

Truth tables are as follows

O F= (A-B)•(3+0) tlooo 10.o), coto) (0:0) (0+1) COD. (1+0) Co.b).(1+1) Ī ToToTo (1.0).(0+0) alesako+1)FA. Bē A B C f . 1o. Tolo Tico.w) @ To lo lo cool T olon 10. Bö 10 i 11 To 10.1 @T (1.000 iloilo 16.03@T 1100C7 TIEDOT@F= AS+BOL - A B Caf: o 10 | 1tops) (0 oo S + C + 001) - 1 + 5 Coⓒo) + Ctool - ot (①1) + (101) at T CI 3 + 1000] - 1+1 (10) +

Some basics.

D = to De A B = AOB AOB - AB

NOT INPUT INPUT OUTPUT OUTPUT А А 0 AND OR XOR INPUT INPUT INPUT OUTPUT OUTPUT OUTPUT A в B DO 0 10 0 101 1 0 1 0 11 0 1 1 1

I have answered all your questions with all necessary basics and details.
Please rate a thumbs up if satisfied.
Have a nice time ahead :)